OpenStudy (anonymous):

The differential equation dy/dx = y/x^2 is has a solution given by:

OpenStudy (anonymous):

a. y = Csqrt(x^2+4) b. y = Ce^(-1/x) c. y=-1/2e^(x^2) + C d. x^2 + y^2 = C e. y=Ce^(-kt)

OpenStudy (anonymous):

The answer is b.

OpenStudy (anonymous):

I will tell you why

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[ \frac{dy}{dx}= \frac {y}{x^2}\\ \frac{dy}{y}= \frac {dx}{x^2}\\ \ln(y)=-\frac {1}x+K\\ y=e^{-\frac {1}x+K}=e^K e^{-\frac {1}x }=C e^{-\frac {1}x } \]

OpenStudy (anonymous):

Thank you, you have been very helpful. :D I can understand your work clearly. Hopefully I past the AP exam next month. Well, it's midnight so I think I will be heading out. Thanks again!

OpenStudy (anonymous):

YW practice some problems on my site http://saab.org